Three particles are fixed on an x axis. Particle 1 of charge q1 is at x = - a, and particle 2 of charge q2 is at x = + a. If their net electrostatic force on particle 3 of charge + Q is to be zero, what must be the ratio q1/q2 when particle 3 is at x = 1.50a? Important: Assume that the charge of particle 1 is positive.
Answer #1

r1 = distance of q1 from q3 = 2.5 a

Magnitude of electric force on q3 by q1 is given as

F1 = k q1 q3/r12

r2 = distance of q2 from q3 = (0.5) a

Magnitude of electric force on q3 by q2 is given as

F2 = k q2 q3/r22

net force on q3 is zero , hence

F1 = F2

k q1 q3/r12 = k q2 q3/r22

q1 /r12 = q2 /r22

q1/q2 = r12/r22

q1/q2 = (2.5 a)2/((0.5) a)2

q1/q2 = 25

Since q1 is positive and q2 is negative, hence the ratio should be negative\

q1/q2 = - 25